OpenStudy (anonymous):

What is the y-intercept of the graph of the equation? 2y + 6 = 5x A. 5/2 (5 over 2) B. -6/5 (Negative 6 over 5) C. 3 D. -3

OpenStudy (anonymous):

It would be 3 correct?

OpenStudy (astrophysics):

You may set it up as y = mx+b form and b will be the y - intercept

OpenStudy (astrophysics):

What do you get when you solve for y?

OpenStudy (anonymous):

Nevermind not 3

pooja195 (pooja195):

It says y intercept just pplug in 0 for x

OpenStudy (astrophysics):

Take it 1 step at a time

pooja195 (pooja195):

*plug

OpenStudy (astrophysics):

Yes! You can also plug in 0 for x and solve for y :)

OpenStudy (anonymous):

I got it, -6/5. Thank you @pooja195 once again! :)

OpenStudy (anonymous):

Sure?

pooja195 (pooja195):

2y + 6 = 5(0) Solve for y your answer is incorrect

OpenStudy (anonymous):

-3.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

2y+6=5x 2y+6=5(0) 2y+6=0 2y=-6 y=-3

pooja195 (pooja195):

Well done :-)

OpenStudy (anonymous):

Thank you guys! Have a nice day! :)